# Visual Odometry with Transformers

### v1

@@ 4.2 Ablation Studies
[…]
VoT delivers real-time prediction. To compare the speed between approaches, we measure their runtime on ScanNet (Dai et al., 2017) scene0000_00 using the same environment. As shown in Fig.˜6, VoT achieves a considerably faster runtime, approximately $3\times$ speedup, compared to existing methods. This efficiency stems from its compact, end-to-end architecture, in contrast to larger 3D models. Moreover, unlike LeapVO, DPVO, and Mast3R-SLAM-VO, VoT does not rely on additional processing steps or bundle adjustment, contributing to both its speed and strong performance.
Limitations and future work. We do not claim VoT can generalize to all videos. Since it is trained on static environments, performance may be limited in dynamic settings. Future improvements could come from scaling to more diverse datasets collected with different devices and careful calibration, as well as expanding coverage to a wider variety of scenes. Employing systematic data curation and incorporating larger models or advanced pre-trained components are also promising directions that could enhance both generalization and overall performance.
5 Conclusion
We presented VoT, an end-to-end method for visual odometry systems based on transformers for direct relative pose prediction. The approach achieves strong results compared to modern state-of-the-art odometry pipelines and large 3D models on the challenging indoor and outdoor datasets. VoT has a flexible architecture easily extensible to various backbones, with competitive results. Importantly, it can generalize to out-of-distribution camera parameters and image data. It exhibits strong scaling behavior, indicating that its performance could be further enhanced with sufficient data and computational resources. Finally, it is considerably faster than the baselines, being more compact than large 3D models, and, unlike other odometry systems, it does not require test-time optimization. We anticipate that future work will further advance end-to-end odometry systems, extending their scalability to a broader range of domains.
Acknowledgements
This work has been financially supported by TomTom, the University of Amsterdam and the allowance of Top consortia for Knowledge and Innovation (TKIs) from the Netherlands Ministry of Economic Affairs and Climate Policy.
Appendix A Appendix
We present further ablation studies and visual examples. Specifically, we test how well VoT scales with different backbone sizes and numbers of input views. We also compare the trajectories predicted by VoT to those from baseline methods, demonstrating the effectiveness of our model.
@@ Appendix A Appendix
Performance of VoT with number of input views. Figure˜7 illustrates how VoT, trained on the full dataset, scales with the number of input views. While the performance plateaus in terms of absolute metrics, it consistently improves in relative metrics. Due to computational constraints, the number of input views in our experiments is limited to a maximum of 12.
@@
Figure 7: Performance of VoT with number of input views on ScanNet (Dai et al., 2017) dataset. As the number of input views increases, absolute translation and rotation errors initially decrease before plateauing. In contrast, relative metrics continue to improve steadily, highlighting the potential of VoT when using a long-range input sequence of views.
@@
Performance of VoT with backbone size. In Fig.˜8, we show how the performance of VoT varies with the size of the backbone. Since CroCO (Weinzaepfel et al., 2022) is only available in a single size, we perform ablation using Dino-V2 (Oquab et al., 2024) backbones. We observe that performance improves as the size of the backbone is increased. Due to computational limits, we cap the backbone size at DinoV2 (Oquab et al., 2024) large.
@@
Figure 8: Performance of VoT with backbone size on ScanNet (Dai et al., 2017) dataset. As the number of backbone parameters increases, the model’s performance improves.
Datasets. ScanNet Dai et al. (2017) is an extensive RGB-D corpus designed for 3D scene understanding in indoor environments, containing over 2 million RGB-D frames from 1,613 unique scenes. Some sequences contain invalid camera poses; in such cases, we use all frames up to the first invalid pose, filtering out 20% of the frames. Additionally, unrealistic jumps in camera trajectories, such as translations spanning several meters between consecutive frames, may occur. To mitigate this, frames with translation magnitudes exceeding 1.5 meters are excluded during training. Our method is evaluated on the entire test split of ScanNet, comprising 98 complete video sequences, using only RGB input. We train our method on KITTI Geiger et al. (2012) by putting 70% of the publicly available data in the training set, and the rest as a hold-out set for testing.
Qualitative Results. In Fig.˜9, we compare several visual odometry methods on representative scenes from the ScanNet (Dai et al., 2017) test set. Each method is evaluated on the entire video sequence using its default settings. All trajectories are visualized without alignment to the ground-truth poses to reflect real-world deployment conditions better. VoT consistently demonstrates accurate and robust camera pose estimation across all evaluated scenes.
LeapVO DPVO CUT3R VGGT Mast3r-SLAM-VO VoT
scene_0717 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
scene_0719 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
scene_0726 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
scene_0744 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
scene_0749 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 9: Qualitative comparisons. Each column shows a method, each row a different scene from the ScanNet (Dai et al., 2017) test set. We evaluate all methods on the whole trajectory without aligning predictions to ground truth, reflecting realistic deployment conditions. LeapVO and DPVO fail to recover scale, while CUT3R and VGGT exhibit significant drift. Mast3r-SLAM-VO estimates poses only at sparse keyframes. VoT consistently achieves robust and accurate reconstructions across diverse scenarios.

### v2

@@ Acknowledgements
Supplementary Material
6 Appendix
We test how well VoT scales with different backbone sizes and numbers of input views.
@@ 6 Appendix
Performance of VoT with number of input views. Figure 6 illustrates how VoT, trained on the full dataset, scales with the number of input views. While the performance plateaus in terms of absolute metrics, it consistently improves in relative metrics. Due to computational constraints, the number of input views in our experiments is limited to a maximum of 12.
@@
Figure 6: Performance of VoT with number of input views on ScanNet [dai2017scannet] dataset. As the number of input views increases, absolute translation and rotation errors initially decrease before plateauing. In contrast, relative metrics continue to improve steadily, highlighting the potential of VoT when using a long-range input sequence of views.
@@
Performance of VoT with backbone size. In Fig. 7, we show how the performance of VoT varies with the size of the backbone. Since CroCO [philippe2022croco] is only available in a single size, we perform ablation using Dino-V2 [oquab2024dinov2] backbones. We observe that performance improves as the size of the backbone is increased. Due to computational limits, we cap the backbone size at DinoV2 [oquab2024dinov2] large.
@@
Figure 7: Performance of VoT with backbone size on ScanNet [dai2017scannet] dataset. As the number of backbone parameters increases, the model’s performance improves.